Find the energy required for Kamryn to vaporize 0.65 kg of gold. (mLv)


1
Expert's answer
2021-06-09T08:05:17-0400

Assume that the initial temperature of gold is "t_0=0\u00b0" .


"Q_1=cm\Delta t=129\cdot0.65\cdot1064=89265\ (J)"


"Q_2=\lambda m=66200\cdot0.65=43030\ (J)"


"Q_3=rm=1736340\cdot0.65=1128621\ (J)"


"Q=Q_1+Q_2+Q_3=89265+43030+1128621=1260916\ (J)" . Answer